OpenStudy (anonymous):

Which is an equation of a circle with center (2, -10) and radius 3?

OpenStudy (im_an_angel):

ok that has something to do with pi * radius squared but i have no idea

OpenStudy (anonymous):

what do you mean @im_an_angel

OpenStudy (anonymous):

anybody know?

OpenStudy (anonymous):

let me solve

OpenStudy (anonymous):

(x 2)2 + (y + 3)2 = 9

OpenStudy (anonymous):

I hope this multiple choice

OpenStudy (anonymous):

(x + 10)2 + (y - 2)2 = 9 (x - 2)2 + (y + 10)2 = 9

OpenStudy (anonymous):

which one?

OpenStudy (anonymous):

second one

OpenStudy (anonymous):

yerp!

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

yea man you welcome

OpenStudy (anonymous):

:))